OpenStudy (anonymous):

Suppose your normal body temperature is 98 and 3/5 degrees Fahrenheit. If your body temperature goes up 2 and 4/5 degrees on Monday and then down 1 and 1/5 degrees on Tuesday, what is your temperature on Tuesday? On Tuesday, your temperature was what degrees. Enter your answer as a reduced mixed number or as a whole number.

OpenStudy (anonymous):

\[98\frac{3}{5}+2\frac{4}{5}-1\frac{1}{5}=?\]

OpenStudy (anonymous):

then what i do

OpenStudy (anonymous):

 you understand why?